Tom traveled 250 miles east of St Louis. For most of the trip he averaged 70mph, but for one period of time he was slowed to 10mph due to a major accident. If the total time of travel was 7 hours, how amny miles did he drive at the reduced speed?
Answer by ewatrrr(24785)

Hi,
Let x be the time at 70mph. Then (7hr-x) would be the time left to traved at 10mph
d = r*t
.
question states
70t + 10(7hr - t) = 250mi
.
simplify and solve
70t + 70hr - 10t = 250mi
60t = 180mi
t = 3 hrs (time driving at 70mph)
.
4hr would be the ime left to be driving at the 10mph reduced speed
.
check your answer
210mi + 40mi = 250 mi